# ELU-GCN: Effectively Label-Utilizing Graph Convolutional Network

### v1

@@ Abstract
The message-passing mechanism of graph convolutional networks (i.e., GCNs) enables label information to be propagated to a broader range of neighbors, thereby increasing the utilization of labels. However, the label information is not always effectively utilized in the traditional GCN framework. To address this issue, we propose a new two-step framework called ELU-GCN. In the first stage, ELU-GCN conducts graph learning to learn a new graph structure (i.e., ELU-graph), which enables GCNs to effectively utilize label information. In the second stage, we design a new graph contrastive learning on the GCN framework for representation learning by exploring the consistency and mutually exclusive information between the learned ELU graph and the original graph. Moreover, we theoretically demonstrate that the proposed method can ensure the generalization ability of GCNs. Extensive experiments validate the superiority of the proposed method.
@@ 1 Introduction
Graph Convolutional Networks (GCNs) (Kipf & Welling, 2017; Gasteiger et al., 2018; Huang et al., 2023a; Xu et al., 2018; Hamilton et al., 2017) have demonstrated remarkable capabilities, primarily due to their ability to propagate label information. This capability has driven their widespread applications in semi-supervised learning. To do this, GCN propagates the representations of unlabeled neighbors to labeled nodes by message passing mechanism, thereby enabling label information to supervise not only the labeled nodes but also their unlabeled neighbors (Ji et al., 2023; Dong et al., 2021). Consequently, the framework of optimizing label utilization in GCNs (LU-GCN) has become an increasingly prominent research topic (Wang et al., 2021; Yue et al., 2022; Yu et al., 2022).
@@
Previous LU-GCN can be partitioned into three categories, i.e., self-training methods, combination methods, and graph learning methods. self-training methods (Dong et al., 2021; Li et al., 2018; Sun et al., 2020; Ji et al., 2023) select unlabeled nodes with the highest classification probability by GCN as training data with pseudo-labels, and thus adding the number of labels to improve the GCN. Combination methods (Wang et al., 2021; Yue et al., 2022; Shi et al., 2021) regard the labels as the augment features so that labels can be used for both representation learning and classification tasks. The feature propagation mechanism allows GCNs to use labels to supervise the representation of both the node itself (i.e., traditional label utilization) and its unlabeled neighbors (i.e., neighboring label utilization). However, the two LU-GCN methods mentioned above primarily focus on optimizing traditional label utilization, neglecting the critical importance of neighboring label utilization in semi-supervised scenarios. Yet, due to noise in the original graph structure, GCNs often struggle to effectively utilize the neighboring labels. To address this issue, recent graph learning methods (Zheng et al., 2020; Luo et al., 2021; Liu et al., 2022) are designed to improve the relationship of every node and its neighbors by updating the graph structure, and thus may potentially improve the neighboring label utilization. For example, Bi et al. (Bi et al., 2022) adopt the own and neighbors’ label similarity to rewire the graph, which can make features propagate on the same category nodes as possible.
@@
Although existing graph learning methods have achieved promising performance, there are still some limitations that need to be addressed. First, previous methods have used heuristic approaches or downstream tasks to learn the graph structure, but they have not explored what kind of graph structures can make GCNs effectively utilize label information. As a result, the graph structures in their methods cannot guarantee that the GCN effectively utilizes the label information. Second, existing graph learning methods fail to explore both the consistency information and the mutually exclusive information between the new graph and the original graph, where they have consistent information (i.e., consistency (Xu et al., 2024)), which helps recognize the node effectively, and every graph contains unique and useful information different from another graph, i.e., mutually exclusive information (Wang et al., 2017).
@@
In this paper, to address the above issues, different from previous structure improvement methods, we investigate a new framework, i.e., Effectively Label-Utilizing GCN (ELU-GCN for brevity), to conduct effective GCN. To achieve this, we first explore the influence of each class provided by labeled nodes on every unlabeled node. We then optimize the graph structure (i.e., ELU-graph) by ensuring that the predictions of GCN align with the primary class information. This ensures that the GCN with the ELU-graph can effectively utilize the label information, thereby addressing challenge (i). Moreover, we address challenge (ii) by designing contrasting constraints to bring the consistency information between two graph views (i.e., the original graph and the ELU-graph) closer and push the mutually exclusive information further apart. Finally, we theoretically analyze that the proposed ELU-graph can not only ensure GCN to effectively utilizes labels, but also improve the generalization ability of the model. Compared with previous methods [^1] Related works are summarized in Appendix C., our main contributions can be summarized as follows:
  • To the best of our knowledge, we are the first attempt to study the limitation of GCNs that cannot effectively utilize labels in the graph framework. Moreover, we provide a quantitative framework to analyze which part of the nodes cannot effectively utilize the label information.
  • We propose to adaptively construct the ELU-graph, which enables the GCN to utilize label information effectively. Furthermore, we design a contrastive loss to leverage the consistency and the mutually exclusive information between the ELU graph and the original graph.
  • We theoretically prove that ELU-graph can ensure the generalization ability of GCN and we experimentally manifest the effectiveness of the proposed method across a variety of datasets, compared with numerous state-of-the-art methods.
@@ 2.1 Motivation
$\mathcal{L}_{\mathrm{DNN}}$ $=\mathrm{CE}(f_{\theta}(\mathbf{X}),\mathbf{Y})=-\sum_{i\in V_{l},k\in C}\nolimits y_{ik}(\log f_{ik})$ (1)
$\mathcal{L}_{\mathrm{GCN}}$ $=\mathrm{CE}(\widehat{\mathbf{A}}f_{\theta}(\mathbf{X}),\mathbf{Y})=-\sum_{i\in V_{l},k\in C}\nolimits y_{ik}(\log\sum_{j\in\mathcal{N}_{i}}\widehat{a}_{ij}f_{jk}),$
where $\theta$ is the parameters of the function $f$. In Eq. (1), the cross entropy loss of DNN is a one-to-one mapping between the feature space and the label space because every label $y_{i}$ ($l=1,...,n$) is only used to supervise the representation learning of one node $v_{i}$. The mapping $f$ efficiently captures the pattern and distribution of labeled nodes, but it overlooks unlabeled nodes so that the generalization ability of unlabeled nodes is limited. In contrast, the cross entropy loss of the GCN is a one-to-many mapping because its message-passing mechanism can propagate the information from labeled nodes to their neighbors including labeled nodes and unlabeled nodes. As a result, every label $y_{i}$ is used to supervise the representation learning of both labeled nodes and unlabeled nodes, as shown in the second row of Eq. (1). Hence, unlabeled nodes in the GCN are able to use the label information of labeled nodes to improve the learning of their representations. Obviously, it is very important to guarantee that unlabeled nodes effectively utilize label information under the GCN framework. However, to the best of our knowledge, no research has focused on this issue. To address this issue, we first quantify the influence of every class on unlabeled nodes, and then make the class with the highest influence (i.e., probability) on unlabeled nodes consistent with the prediction of the GCN to effectively utilize the label information.
The recent study in (Xu et al., 2018) reveals that nodes follow the way of random walks to affect other nodes on the graph. Therefore, in this paper, we extend it to obtain the influence of every class of labeled nodes to the unlabeled node by Theorem 2.1, whose proof is provided in Appendix B.1.
Theorem 2.1.
@@
Based on Theorem 2.1, LPA can be utilized to calculate the probability of every class for unlabeled nodes in the GCN framework. The class with the highest probability is considered the most important for the unlabeled node, as it contributes the most label information. If the class that carries the most label information to a node in the GCN framework is the same as the GCN prediction, it indicates that the label information propagated by the message passing mechanism of the GCN positively influences the classification result for that node. In this way, this node is regarded as effectively utilizing the label information. We provide a case study to illustrate this in Figure 1 and give a formal definition as follows.
@@
(Effective label-utilization) The GCN effectively utilizes label information if the prediction of GCN is consistent with the output of LPA, i.e.,
@@
where $V_{\mathrm{ELU}}$ and $V_{\mathrm{NELU}}$ (i.e., $V_{\mathrm{NELU}}=\{V|\mathrm{LPA}(\mathcal{G})\neq\mathrm{GCN}(\mathcal{G})\}$), respectively, represent the node set which effectively utilizes the label information and the node set which does not effectively utilizes the label information in the GCN framework.
In real applications, not all unlabeled nodes in GCN frameworks may effectively utilize the label information due to all kinds of reasons, including noise and the distribution of labeled nodes in the graph. Figure 2 shows that not all nodes effectively use label information in the GCN framework (i.e., Figure 2 (a)) and the classification accuracy of $V_{\mathrm{NELU}}$ is lower than that of $V_{\mathrm{ELU}}$ in the same datasets (i.e., Figure 2 (b)). Obviously, NELU nodes influence the effectiveness of the GCN. To address this issue, first, it is crucial to make unlabeled nodes effectively utilize label information. Since label information is propagated through the graph structure. As a result, the graph structure will be updated. Second, the original graph structures often contain noise to influence the message-passing mechanism. Hence, graph learning is obviously a feasible solution.
@@ 2.2 ELU Graph
Previous graph learning methods generally use either heuristic methods or downstream tasks to conduct graph learning, i.e., updating the graph structure. For example, Pro-GNN (Jin et al., 2020) updates the graph structure through a heuristic approach to constrain the sparsity and smoothness of the graph. PTD-Net (Luo et al., 2021) updates the graph structure by the downstream task, such as the node classification task. However, heuristic methods rely on predefined rules, making it difficult for unlabeled nodes to fully access label-related global information. Downstream task methods focus too much on the performance of labeled nodes, neglecting the role of unlabeled nodes in the graph structure. Therefore, these efforts cannot ensure unlabeled nodes effectively utilize label information. To solve this issue, based on Definition 2.2, we investigate new graph learning methods that ensure unlabeled nodes effectively utilize label information.
Specifically, denoting the adjacency matrix $\mathbf{S}$ as the ELU graph can ensure the GCN effectively uses the label information, we use Theorem 2.1 to measure the influence of each class on every unlabeled node by the LPA:
@@
In Eq. (5), the prediction of GCN and the influence of each class are encouraged to be consistent for every node. This item can make all nodes satisfy $\mathrm{LPA}(\mathcal{G})=\mathrm{GCN}(\mathcal{G})$ in Eq. (2), i.e., this objective function can ensure all nodes can effectively utilize label information by GCN. Therefore, we can obtain the $\mathbf{S}$ through the optimization algorithm by minimizing the Eq. (5). However, there are some problems with the above objective function. First, it is impracticable to solve the above problem directly, as it has a trivial solution: $s_{i,j}=0,\forall i,\forall j$. Second, LPA generates the prediction for every labeled node to possibly revise the original labels, i.e., the ground truth, adding noisy labels for representation learning. To overcome the above issues, We propose to iteratively update in two steps, i.e., update labels by LPA and update the graph structure $\mathbf{S}$.
In the first step, we calculate the result of LPA $\mathbf{Q}^{(i)}$, i.e., $\mathbf{Q}^{(i)}=\mathbf{S}^{(i-1)}\mathbf{Q}^{(i-1)}$, ($i=1,\dots,k$), where $\mathbf{Q}^{(0)}=\mathbf{Y}$ and we initialize $\mathbf{S}^{(0)}=\mathbf{I_{N}}$. As a result, Eq.(5) is changed as follows:
@@
In the second step, we can obtain its closed-form solution, which is listed as follows and its details are in Appendix B.2:
$\mathbf{S}^{(i)}=\mathbf{H}(\mathbf{Q}^{(i)})^{T}\left(\mathbf{H}\mathbf{H}^{T}+\beta\mathbf{I}_{N}\right)^{-1},$ (7)
@@
$\mathbf{Q}^{(i)}=\mathbf{H}(\mathbf{Q}^{(i-1)})^{T}\left(\frac{1}{\beta}\mathbf{I}_{N}-\frac{1}{\beta^{2}}\mathbf{H}\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}\mathbf{H}^{T}\right)\mathbf{Q}^{(i-1)},s.t.~{}~{}\mathbf{Q}^{(i)}_{l}=\mathbf{Y}_{l},$ (8)
@@
$\mathbf{S}^{*}=\mathbf{H}(\frac{1}{\beta}(\mathbf{Q}^{(k)})^{T}-\frac{1}{\beta^{2}}(\mathbf{Q}^{(k)})^{T}\mathbf{H}\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}\mathbf{H}^{T}).$ (9)
The details of Eq. (9) are listed in Appendix A.2. The pseudocode of calculating Eq. (LABEL:eq:10) and $\mathbf{S}^{*}$ is presented in Algorithm 2. In the implementation, we make $\mathbf{S}^{*}$ sparse by assigning its element less than a threshold as zero, for achieving efficiency. We also use the pseudo labels of ELU nodes to expand the initial $\mathbf{Y}$, for avoiding the issue of limited labels in semi-supervised learning.

### v3

# Enhancing the Influence of Labels on Unlabeled Nodes in Graph Convolutional Networks

@@ Abstract
The message-passing mechanism of graph convolutional networks (i.e., GCNs) enables label information to reach more unlabeled neighbors, thereby increasing the utilization of labels. However, the additional label information does not always contribute positively to the GCN. To address this issue, we propose a new two-step framework called ELU-GCN. In the first stage, ELU-GCN conducts graph learning to learn a new graph structure (i.e., ELU-graph), which allows the additional label information to positively influence the predictions of GCN. In the second stage, we design a new graph contrastive learning on the GCN framework for representation learning by exploring the consistency and mutually exclusive information between the learned ELU graph and the original graph. Moreover, we theoretically demonstrate that the proposed method can ensure the generalization ability of GCNs. Extensive experiments validate the superiority of our method.
Graph Neural Networks, Semi-Supervised Learning, Graph Learning
@@ 1 Introduction
Graph Convolutional Networks (GCNs) (Kipf & Welling, 2017; Gasteiger et al., 2018; Huang et al., 2023a; Xu et al., 2018; Hamilton et al., 2017) have demonstrated remarkable capabilities, primarily due to their ability to propagate label information. This capability has driven their widespread applications in semi-supervised learning. To do this, GCN propagates the representations of unlabeled neighbors to labeled nodes by the message passing mechanism, thereby enabling label information to supervise not only the labeled nodes but also their unlabeled neighbors (Ji et al., 2023; Dong et al., 2021). Consequently, the framework of optimizing label utilization in GCNs (LU-GCN) has become an increasingly prominent research topic (Wang et al., 2021; Yue et al., 2022; Huang et al., 2024; Yu et al., 2022).
@@
Previous LU-GCN can be partitioned into three categories, i.e., self-training methods, combination methods, and graph learning methods. self-training methods (Dong et al., 2021; Li et al., 2018; Sun et al., 2020; Ji et al., 2023) select unlabeled nodes with the highest classification probability by GCN as training data with pseudo-labels, thus adding the number of labels to improve the GCN. Combination methods (Wang et al., 2021; Yue et al., 2022; Shi et al., 2021) regard the labels as the augmented features so that labels can be used for both representation learning and classification tasks. The feature propagation mechanism allows GCNs to use labels to supervise the representation of both the node itself (i.e., traditional label utilization) and its unlabeled neighbors (i.e., neighboring label utilization). However, the two LU-GCN methods mentioned above primarily focus on optimizing traditional label utilization, neglecting the critical importance of neighboring label utilization in semi-supervised scenarios. Yet, due to noise in the original graph structure, GCNs often struggle to utilize the neighboring labels effectively. To address this issue, recent graph learning methods (Zheng et al., 2020; Luo et al., 2021; Liu et al., 2022) are designed to improve the relationship of every node and its neighbors by updating the graph structure, and thus may potentially improve the neighboring label utilization. For example, Bi et al. (2022) adopt the own and neighbors’ label similarity to rewire the graph, which can make features propagate on the same category nodes as possible.
@@
Although existing graph learning methods have achieved promising performance, there are still some limitations that need to be addressed. First, previous methods have used heuristic approaches to learn the graph structure, they have not explored what kind of graph structures can make GCNs effectively propagate the labeled nodes’ information to unlabeled nodes. As a result, on the learned graph structure, the label information propagated via message passing may not always contribute positively to the predictions of GCNs. Second, existing graph learning methods fail to explore both the consistency information and the mutually exclusive information between the new graph and the original graph, where they have consistent information (i.e., consistency (Xu et al., 2024)), which helps recognize the node effectively, and every graph contains unique and useful information different from another graph, i.e., mutually exclusive information (Wang et al., 2017).
@@
In this paper, to address the above issues, different from previous structure improvement methods, we investigate a new framework, i.e., Effectively Label-Utilizing GCN (ELU-GCN for brevity), to conduct effective GCN. To achieve this, we first analyze the influence of each class provided by labeled nodes on every unlabeled node. We then optimize the graph structure (i.e., the ELU-graph) by encouraging the label information propagated through the graph to have a positive impact on GCN’s prediction results. This ensures that the GCN with the ELU-graph can effectively utilize the label information, thereby addressing challenge (i). Moreover, we address challenge (ii) by designing contrasting constraints to bring the consistent information between two graph views (i.e., the original graph and the ELU-graph) closer and push the mutually exclusive information further apart. Finally, we theoretically analyze that the proposed ELU-graph can ensure GCN effectively utilizes labels and improve the generalization ability of the model. Compared with previous methods, our main contributions can be summarized as follows:
  • To the best of our knowledge, this is the first work to investigate the limitation of GCNs in effectively utilizing label information under the graph-based learning paradigm. Moreover, we introduce a quantitative framework to analyze which unlabeled nodes struggle to benefit from label supervision by neighborhood.
  • We propose an adaptive and parameter-free construction of the ELU-graph to enhance GCNs’ ability to utilize label information, particularly for unlabeled nodes. Furthermore, we design a contrastive loss to exploit both the consistency and complementary differences between the ELU-graph and the original graph.
  • We theoretically prove that the ELU-graph construction provides generalization guarantees for GCNs. Empirical results on diverse benchmark datasets further validate the superior performance of our method over competitive baselines.
2 Related Works
This section briefly reviews the topics related to this work, including GCNs and LPA as well as graph structure learning.
2.1 GCNs and LPA
GCNs are the most popular and commonly used model in the field of graph deep learning. Early work attempted to apply the successful convolutional neural network (CNN) to graph structures. For example, CheybNet (Defferrard et al., 2016) first proposes to transform the graph signal from the spatial domain to the spectral domain through the discrete Fourier transform, and then use polynomials to fit the filter shape (i.e., convolution). CheybNet laid the foundation for the development of spectral-domain graph neural networks. The popular GCN was proposed by Kipf et al. (Kipf & Welling, 2017), which is a simplified version of ChebyNet and has demonstrated strong efficiency and effectiveness, thereby promoting the development of the graph deep learning field.
Recently, some works have focused on the combination of LPA and GCN. This is because LPA can characterize the distribution of labels spread on the graph, which can help GCN obtain more category information. Existing combined LPA and GCN methods can be classified into two categories, i.e., pseudo label methods and masked label methods. Pseudo-label methods let the output of LPA serve as the pseudo-labels to guide the representation learning. For example, PTA (Dong et al., 2021) first propagates the known labels along the graph to generate pseudo-labels for the unlabeled nodes, and second, trains normal neural network classifiers on the augmented pseudo-labeled data. GPL (Wu et al., 2024) uses the output of LPA to preserve the edges between nodes of the same class, thereby reducing the intra-class distance. The masked label methods employ LPA as regularization to assist the GCN update parameters or structures. For example, UniMP (Shi et al., 2021) makes some percentage of input label information masked at random, and then predicts it for updating parameters. GCN-LPA (Wang & Leskovec, 2021) also randomly masked a part of the labels and utilized the remaining label nodes to predict them in learning proper edge weights within labeled nodes. Although the above methods achieve excellent results on various tasks, they all overlook whether the additional label information propagated through message passing is effectively utilized by GCNs to improve unlabeled nodes.
2.2 Graph Structure Learning
Graph structure learning is an important technology in the graph field. It can improve the graph structure and infer new relationships between samples, thereby promoting the development of graph representation learning or other fields. Existing Graph structure learning methods can be classified into three categories, i.e., traditional unsupervised graph structure learning methods, supervised graph structure learning methods, and graph rewiring methods.
Traditional unsupervised graph structure learning methods aim to directly learn a graph structure from a set of data points in an unsupervised manner. Early works (Wang & Zhang, 2006; Daitch et al., 2009) exploit the neighborhood information of each data point for graph construction by assuming that each data point can be optimally reconstructed using a linear combination of its neighbors (i.e., $\min_{A}\left\|\mathbf{AX}-\mathbf{X}\right\|_{F}^{2}$). Similarly, (Daitch et al., 2009) introduce the weight (i.e., $\min\sum_{i}\left\|\mathbf{D}_{i,i}\mathbf{X}_{i}-\sum_{j}\mathbf{A}_{i,j}\mathbf{X}_{j}\right\|^{2}$). Smoothness (Jiang et al., 2019) is another widely adopted assumption on natural graph signals; the smoothness of the graph signals is usually measured by the Dirichlet energy (i.e., $\min_{\mathbf{A}}\frac{1}{2}\sum_{i,j}\mathbf{A}_{i,j}\left\|\mathbf{X}_{i}-\mathbf{X}_{j}\right\|^{2}$ and $\min_{\mathbf{L}}\operatorname{tr}\left(\mathbf{X}^{\top}\mathbf{L}\mathbf{X}\right)$). The above objective has inspired many advanced graph structure learning methods.
Supervised graph structure learning methods aim to use the downstream task to supervise the structure learning, which can learn a suitable structure for the downstream task. For example, NeuralSparse (Zheng et al., 2020) and PTDNet (Luo et al., 2021) directly use the adjacency matrix of the graph as a parameter and update the adjacency matrix through the downstream task. SA-SGC (Huang et al., 2023b) learns a binary classifier by distinguishing the edges connecting nodes with the same label and the edges connecting nodes with different labels in the training set, thereby deleting the edges between nodes belonging to different categories in the test set. BAGCN (Zhang et al., 2024) uses metric learning to obtain new graph structures and learns suitable metric spaces through downstream tasks.
The goal of graph rewiring methods is to prevent the over-squashing (Alon & Yahav, 2021) problem. For example, FA (Alon & Yahav, 2021) proposed to use a fully connected graph as the last layer of GCN to overcome over-squashing. SDRF (Topping et al., 2022), SJLR (Giraldo et al., 2023), and BORF (Nguyen et al., 2023) aim to enhance the curvature of the neighborhood by rewiring connecting edges with small curvature. They increase local connectivity in the graph topology, indirectly expanding the influence range of labels. Despite their success, existing graph structure learning methods fail to guarantee that the learned structures enable GNNs to propagate label information to unlabeled nodes effectively.
@@ 3.1 Motivation
$\mathcal{L}_{\mathrm{DNN}}$ $=\mathrm{CE}(f_{\theta}(\mathbf{X}),\mathbf{Y})=-\sum_{i\in V_{l},k\in C}y_{ik}(\log f_{ik})$ (1)
$\mathcal{L}_{\mathrm{GCN}}$ $=\mathrm{CE}(\widehat{\mathbf{A}}f_{\theta}(\mathbf{X}),\mathbf{Y})=-\sum_{i\in V_{l},k\in C}y_{ik}(\log\sum_{j\in\mathcal{N}_{i}}\widehat{a}_{ij}f_{jk}),$
where $\theta$ is the parameters of the function $f$. In Eq. (1), the cross entropy loss of DNN is a one-to-one mapping between the feature space and the label space because every label $y_{i}$ ($l=1,...,n$) is only used to supervise the representation learning of one node $v_{i}$. The mapping $f$ efficiently captures the pattern and distribution of labeled nodes, but it overlooks unlabeled nodes so that the generalization ability of unlabeled nodes is limited. In contrast, the cross entropy loss of the GCN is a one-to-many mapping because its message-passing mechanism can propagate the information from labeled nodes to their neighbors including labeled nodes and unlabeled nodes. As a result, every label $y_{i}$ is used to supervise the representation learning of both labeled nodes and unlabeled nodes, as shown in the second row of Eq. (1). Hence, unlabeled nodes in the GCN are able to use the label information of labeled nodes to improve the learning of their representations. Obviously, it is critical to ensure that the label information propagated to unlabeled nodes makes a positive contribution to GCN. However, to the best of our knowledge, this issue has not yet received sufficient attention. A related work, GCN-LPA (Wang & Leskovec, 2021), investigates the influence of labeled nodes; however, it primarily focuses on reinforcing mutual influence among labeled nodes, while overlooking their impact on unlabeled nodes. To address this issue, we first quantify how much influence each class, through its labeled nodes, has on the unlabeled nodes.
The recent study in (Xu et al., 2018) reveals that nodes follow the way of random walks to affect other nodes on the graph. Therefore, in this paper, we extend it to obtain the influence of every class of labeled nodes on the unlabeled node by Proposition 3.1, whose proof is provided in Appendix B.1.
Proposition 3.1.
@@
Based on Proposition 3.1, LPA can be used to estimate the class-wise probability for unlabeled nodes in the GCN framework. The class with the highest probability is regarded as the most influential, as it contributes the most label information to the node. If this most influential class matches the GCN prediction, it indicates that the label information provided through message passing positively influences the GCN’s classification for that node. In this way, the GCN is regarded as having effectively utilized the label information on this node. We provide a case study to illustrate this in Figure 1 and give a formal definition as follows.
@@
(Effective label-utilization) The GCN effectively utilizes label information if the prediction of the GCN is consistent with the output of LPA. This condition is defined at the node level:
@@
where $V_{\mathrm{ELU}}$ and $V_{\mathrm{NELU}}$ (i.e., $V_{\mathrm{NELU}}=\{V|\mathrm{LPA}(\mathcal{G})\neq\mathrm{GCN}(\mathcal{G})\}$), respectively, represent the sets of nodes on which the GCN effectively and ineffectively utilizes label information, respectively.
In real applications, not all unlabeled nodes in GCN frameworks may effectively utilize the label information due to all kinds of reasons, including noise and the distribution of labeled nodes in the graph. Figure 2 shows that not all unlabeled nodes effectively use label information in the GCN framework (i.e., Figure 2 (a)) and the classification accuracy of $V_{\mathrm{NELU}}$ is lower than that of $V_{\mathrm{ELU}}$ in the same datasets (i.e., Figure 2 (b)). Obviously, NELU nodes influence the effectiveness of the GCN. To address this issue, first, it is crucial to make unlabeled nodes effectively utilize label information. Since label information is propagated through the graph structure. As a result, the graph structure will be updated. Second, the original graph structures often contain noise to influence the message-passing mechanism. Hence, graph learning is obviously a feasible solution.
@@ 3.2 ELU Graph
Previous graph learning methods generally use either heuristic methods or downstream tasks to conduct graph learning, i.e., updating the graph structure. For example, Pro-GNN (Jin et al., 2020) updates the graph structure through a heuristic approach to constrain the sparsity and smoothness of the graph. PTD-Net (Luo et al., 2021) updates the graph structure by the downstream task, such as the node classification task. However, heuristic methods rely on predefined rules, making it difficult for unlabeled nodes to fully access label-related global information. Downstream task methods focus too much on the performance of labeled nodes, neglecting the role of unlabeled nodes in the graph structure. Therefore, these methods fail to guarantee that the GCN effectively leverages label information for unlabeled nodes. To solve this issue, based on Definition 3.2, we investigate new graph learning methods that ensure unlabeled nodes effectively utilize label information.
Specifically, denoting the adjacency matrix $\mathbf{S}$ as the ELU graph can ensure the GCN effectively uses the label information. We use Proposition 3.1 to measure the influence of each class on every unlabeled node by the LPA:
@@
In Eq. (5), the prediction of GCN and the influence of each class are encouraged to be consistent for every node. This item can make all nodes satisfy $\mathrm{LPA}(\mathcal{G})=\mathrm{GCN}(\mathcal{G})$ in Eq. (2), i.e., this objective function can ensure all nodes can effectively utilize label information by GCN. Therefore, we can obtain the $\mathbf{S}$ through the optimization algorithm by minimizing the Eq. (5). However, there are some problems with the above objective function. First, it is impractical to solve the above problem directly, as it has a trivial solution: $s_{i,j}=0,\forall i,\forall j$. Second, LPA generates the prediction for every labeled node to possibly revise the original labels, i.e., the ground truth, adding noisy labels for representation learning. To overcome the above issues, We propose to iteratively update in two steps, i.e., update labels by LPA and update the graph structure $\mathbf{S}$.
In the first step, we calculate the result of LPA $\mathbf{Q}^{(i)}$, i.e., $\mathbf{Q}^{(i)}=\mathbf{S}^{(i-1)}\mathbf{Q}^{(i-1)}$, ($i=1,\dots,k$), where $\mathbf{Q}^{(0)}=\mathbf{Y}$. As a result, Eq.(5) is changed as follows:
@@
In the second step, we can obtain its closed-form solution as follows:
where $Tr(\cdot)$ indicates the trace of matrix. Then we have
$\frac{\partial\mathcal{L}}{\partial\mathbf{S}}=-2\mathbf{Q}^{(i)}\mathbf{H}^{T}+2\mathbf{S}\mathbf{H}\mathbf{H}^{T}+2\beta\mathbf{S}$ (8)
Let Eq. (8) equal to 0, we can obtain the closed-form solution $\mathbf{S}^{(i)}$ i.e.,
$\mathbf{S}^{(i)}=\mathbf{Q}^{(i)}\mathbf{H}^{T}\left(\mathbf{H}\mathbf{H}^{T}+\beta\mathbf{I}_{N}\right)^{-1}.$ (9)
@@
$\mathbf{Q}^{(i)}=$ (10)
$\mathbf{Q}^{(i-1)}\mathbf{H}^{T}\left(\frac{1}{\beta}\mathbf{I}_{N}-\frac{1}{\beta^{2}}\mathbf{H}\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}\mathbf{H}^{T}\right)\mathbf{Q}^{(i-1)},$
$s.t.~{}~{}\mathbf{Q}^{(i)}_{l}=\mathbf{Y}_{l},$
@@
$\mathbf{S}^{*}=\mathbf{Q}^{(k)}(\frac{1}{\beta}\mathbf{H}^{T}-\frac{1}{\beta^{2}}\mathbf{H}^{T}\mathbf{H}\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}\mathbf{H}^{T}).$ (11)
The details of Eq. (11) are listed in Section A.2. The pseudocode of calculating Eq. (LABEL:eq:10) and $\mathbf{S}^{*}$ is presented in Algorithm 1. In the implementation, we make $\mathbf{S}^{*}$ sparse by assigning its element less than a threshold as zero, for achieving efficiency. We also use the pseudo labels of ELU nodes to expand the initial $\mathbf{Y}$, for avoiding the issue of limited labels in semi-supervised learning.